# MIRe: Enhancing Multimodal Queries Representation via Fusion-Free Modality Interaction for Multimodal Retrieval

### v2

@@ Abstract
Recent multimodal retrieval methods have endowed text-based retrievers with multimodal capabilities by utilizing pre-training strategies for visual-text alignment. They often directly fuse the two modalities for cross-reference during the alignment to understand multimodal queries. However, existing methods often overlook crucial visual information due to a text-dominant issue, which overly depends on text-driven signals. In this paper, we introduce MIRe, a retrieval framework that achieves modality interaction without fusing textual features during the alignment. Our method allows the textual query to attend to visual embeddings while not feeding text-driven signals back into the visual representations. Additionally, we construct a pre-training dataset for multimodal query retrieval by transforming concise question-answer pairs into extended passages. Our experiments demonstrate that our pre-training strategy significantly enhances the understanding of multimodal queries, resulting in strong performance across four multimodal retrieval benchmarks under zero-shot settings. Our code is publicly available: https://github.com/yeongjoonJu/MIRe.
@@ 1 Introduction
In this paper, we introduce MIRe, a retrieval framework that achieves modality interaction without fusing textual features during the alignment stage. Instead of directly merging both modalities, MIRe allows the textual query to attend to patch-level visual embeddings without feeding text-driven signals back into the visual representations. We then fuse the two modalities during the relevance scoring stage based on a late-interaction mechanism Khattab and Zaharia (2020). This design alleviates the dependency on text-driven signals in the context of knowledge retrieval using a multimodal query. Furthermore, we construct a pre-training dataset by transforming multimodal query-response pairs into extensive passages via our response-to-passage conversion process that utilizes solely a text retrieval model. The constructed dataset requires the integration of both modalities to match a desired passage during training, enabling the model to link image understanding with complex textual queries. Our experiments demonstrate that our pre-training strategy significantly enhances multimodal query understanding for knowledge retrieval, resulting in strong performance across four multimodal retrieval benchmarks under zero-shot settings.
@@ 2 Related Work
The transition from traditional text queries to multimodal queries has marked a significant evolution in information retrieval Luo et al. (2021a). Early methods focused on converting images into textual representations, such as captions Qu et al. (2021); Gao et al. (2022) and object tags Gui et al. (2022); Yang et al. (2022). EnFoRe Wu and Mooney (2022) and DEDR Salemi et al. (2023) improve image-query representations derived from a multimodal encoder with generated entities and captions, respectively. Whereas these approaches utilize DPR Karpukhin et al. (2020) based on a single embedding for retrieval, FLMR Lin et al. (2023) refines multimodal queries by incorporating RoIs and generated captions with the late-interaction mechanism. ReViz Luo et al. (2023) represents an end-to-end multimodal retrieval system that removes the dependency on intermediate modules by pre-training on the VL-ICT, which automatically constructs a pre-training dataset by applying the Inverse Cloze Task (ICT) Lee et al. (2019) to a multimodal knowledge base. UniIR Wei et al. (2024) proposes an instruction-guided multimodal retriever along with its benchmark. They design two variants of the model architecture for modality interaction: score-level fusion and feature-level fusion based on CLIP and BLIP Li et al. (2022). VISTA Zhou et al. (2024a) introduces an in-depth fusion strategy by prepending visual tokens to the input of a text retrieval model to enhance multimodal understanding. PreFLMR Lin et al. (2024) extends FLMR to investigate the scalability of multimodal retrievers under the late-interaction mechanism. In contrast to previous methods that rely heavily on text information within multimodal queries, we address the text-dominant issue in multimodal query representations caused by the direct intervention of textual features. We also adopt the late-interaction mechanism to fuse modalities during the scoring stage.
@@ 3.3 Model Architecture
Query-guided Attentive Pooling. Our architecture aims to achieve modality interaction without directly incorporating textual features in the pre-training stage for multimodal alignment. To this end, we introduce a query-guided attentive pooling module. This module retrieves visual information required by the textual query $T$ from $V_{m}\in\mathbb{R}^{l_{v}\times d_{v}}$ and then aggregates the visual information based on its relevance to tokens within the textual query, where $l_{v}$ denotes the number of image patches. Employing $E_{t}$ as query vectors, attention scores $\mathcal{A}\in\mathbb{R}^{h\times l_{t}\times l_{v}}$ are calculated as follows:
@@
where $\mathcal{V}_{m}$ is produced via operations identical with $\mathcal{K}_{m}$. Unlike the standard cross-attention mechanism, we apply mean-pooling along the sequence dimension without a residual connection, yielding $h$ visual embeddings. In this modality interaction process, we only leverage textual embeddings $E_{t}$ to calculate $\mathcal{A}$ as relevance scores for $T$ without direct fusion.
@@ 4 Experiments
Table 1: Zero-shot performance of MIRe and comparison methods. Note that FLMR was only pre-trained on the WiT dataset. PreFLMR were trained using our dataset and experimental settings. Bold indicates the highest performance, while underline signifies the second highest performance.
@@ 4.1 Setup
Comparison Methods. We benchmark our MIRe model against a diverse set of baseline models that employ pre-training stages for visual-text alignment: CLIP Radford et al. (2021), FLMR Lin et al. (2023), ReViz Luo et al. (2023), PreFLMR Lin et al. (2024), and VISTA Zhou et al. (2024a). Both FLMR and PreFLMR utilize the same vision and text encoders as our model, where FLMR was pre-trained with a subset of the WiT dataset. For direct comparison, PreFLMR was trained using the same pre-training procedure as our model, thereby highlighting the distinct advantages of our model architecture.
@@ 5 Discussion
Text-dominant Issue. In Fig. 5(a), both PreFLMR and w/ Residual exhibit faster loss convergence compared to our model, suggesting that directly leveraging text features accelerates optimization. However, as shown in Fig. 5(b), the accelerated convergence does not translate to improved performance, with PreFLMR and w/ Residual underperforming relative to our model. The text-dominant issue is further exacerbated when using VL-ICT, a dataset that constructs pseudo-queries from passages with an image, reinforcing the reliance on textual features. Such behavior highlights the text-dominant issue, where excessive dependence on text features during alignment hinders the model’s ability to fully leverage multimodal information. Fig. 6 illustrates this effect by visualizing the alignment of multimodal query embeddings $Q$ with passage embeddings $D$. In (a), when $E_{t}$ is explicitly used during alignment $Q$ embeddings (orange) remain largely separated from the passage space, indicating poor alignment. In contrast, (b) demonstrates that our method effectively incorporates $Q$ embeddings (green) into the linguistic space, improving alignment. These results suggest that excessive reliance on text features inhibits the multimodal query embeddings from adapting properly to the passage space, reinforcing the text-dominant issue observed in retrieval performance.

### v3

@@ Abstract
Recent multimodal retrieval methods have endowed text-based retrievers with multimodal capabilities by utilizing pre-training strategies for visual-text alignment. They often directly fuse the two modalities for cross-reference during the alignment to understand multimodal queries. However, existing methods often overlook crucial visual information due to a text-dominant issue, which overly depends on text-driven signals. In this paper, we introduce MIRe, a retrieval framework that achieves modality interaction without fusing textual features during the alignment. Our method allows the textual query to attend to visual embeddings while not feeding text-driven signals back into the visual representations. Additionally, we construct a pre-training dataset for multimodal query retrieval by transforming concise question-answer pairs into extended passages. Our experiments demonstrate that our pre-training strategy significantly enhances the understanding of multimodal queries, resulting in strong performance across four multimodal retrieval benchmarks under zero-shot settings. Moreover, our ablation studies and analyses explicitly verify the effectiveness of our framework in mitigating the text-dominant issue. Our code is publicly available: https://github.com/yeongjoonJu/MIRe
@@ 1 Introduction
To address these issues, we introduce MIRe, a retrieval framework that achieves modality interaction without fusing textual features during the alignment stage. Instead of directly merging both modalities, MIRe allows the textual query to attend to patch-level visual embeddings without feeding text-driven signals back into the visual representations. We then fuse the two modalities during the relevance scoring stage based on a late-interaction mechanism Khattab and Zaharia (2020). This design alleviates the dependency on text-driven signals in the context of knowledge retrieval using a multimodal query. Furthermore, we construct a pre-training dataset by transforming multimodal query-response pairs into extensive passages via our response-to-passage conversion process that utilizes solely a text retrieval model. The constructed dataset requires the integration of both modalities to match a desired passage during training, enabling the model to link image understanding with complex textual queries. Our experiments demonstrate that our pre-training strategy significantly enhances multimodal query understanding for knowledge retrieval, resulting in strong performance across four multimodal retrieval benchmarks under zero-shot settings.
@@ 2 Related Work
The transition from traditional text queries to multimodal queries has marked a significant evolution in information retrieval Luo et al. (2021a). Early methods focused on converting images into textual representations, such as captions Qu et al. (2021); Gao et al. (2022) and object tags Gui et al. (2022); Yang et al. (2022). EnFoRe Wu and Mooney (2022) and DEDR Salemi et al. (2023) improve image-query representations derived from a multimodal encoder with generated entities and captions, respectively. The OVEN dataset Hu et al. (2023) has also provided insights into multimodal entity recognition. Whereas most of these approaches utilize DPR Karpukhin et al. (2020) based on a single embedding for retrieval, FLMR Lin et al. (2023) refines multimodal queries by incorporating RoIs and generated captions with the late-interaction mechanism. ReViz Luo et al. (2023) represents an end-to-end multimodal retrieval system that removes the dependency on intermediate modules by pre-training on the VL-ICT, which automatically constructs a pre-training dataset by applying the Inverse Cloze Task (ICT) Lee et al. (2019) to a multimodal knowledge base. UniIR Wei et al. (2024) proposes an instruction-guided multimodal retriever along with its benchmark. They design two variants of the model architecture for modality interaction: score-level fusion and feature-level fusion based on CLIP and BLIP Li et al. (2022). VISTA Zhou et al. (2024a) introduces an in-depth fusion strategy by prepending visual tokens to the input of a text retrieval model to enhance multimodal understanding. PreFLMR Lin et al. (2024) extends FLMR to investigate the scalability of multimodal retrievers under the late-interaction mechanism. In contrast to previous methods that rely heavily on text information within multimodal queries, we address the text-dominant issue in multimodal query representations caused by the direct intervention of textual features. We also adopt the late-interaction mechanism to fuse modalities during the scoring stage.
@@ 3.3 Model Architecture
Query-guided Attentive Pooling. Our architecture aims to achieve modality interaction without directly incorporating textual features in the pre-training stage for multimodal alignment, thereby mitigating the text dominance issue. To this end, we introduce a query-guided attentive pooling module employing an attention layer. This module retrieves visual information required by the textual query $T$ from $V_{m}\in\mathbb{R}^{l_{v}\times d_{v}}$ and then aggregates the visual information based on its relevance to tokens within the textual query, where $l_{v}$ denotes the number of image patches. Employing $E_{t}$ as query vectors, attention scores $\mathcal{A}\in\mathbb{R}^{h\times l_{t}\times l_{v}}$ are calculated as follows:
@@
where $\mathcal{V}_{m}$ is produced via operations identical with $\mathcal{K}_{m}$. Unlike the standard cross-attention mechanism, we apply mean-pooling along the sequence dimension without a residual connection, yielding $h$ visual embeddings. In this modality interaction process, we only leverage textual embeddings $E_{t}$ to calculate $\mathcal{A}$ as relevance scores for $T$ without direct fusion. Our module alleviates the text dominance issue by explicitly limiting textual information during visual representation alignment, as demonstrated in our empirical analyses (see Sec. 5).
@@ 4 Experiments
Table 1: Zero-shot performance of MIRe and comparison methods. Note that FLMR was only pre-trained on the WiT dataset. PreFLMR were trained using our dataset and experimental settings. Bold indicates the highest performance.
@@ 4.1 Setup
Comparison Methods. We benchmark our MIRe model against a diverse set of baseline models that employ pre-training stages for visual-text alignment: CLIP Radford et al. (2021), FLMR Lin et al. (2023), ReViz Luo et al. (2023), PreFLMR Lin et al. (2024), and VISTA Zhou et al. (2024a). UniIR Wei et al. (2024), which requires an explicit instruction input, is also included; we follow their protocol and use the instruction "Retrieve a passage that answers the given query about the image" during evaluation. Both FLMR and PreFLMR utilize the same vision and text encoders as our model, where FLMR was pre-trained with a subset of the WiT dataset. For direct comparison, PreFLMR was trained using the same pre-training procedure as our model, thereby highlighting the distinct advantages of our model architecture. For zero-shot (ZS) evaluation, we exclude baselines requiring external supervision for fairness, while for few-shot (FS) evaluation, models utilizing supervision datasets are included.
@@ 5 Discussion
Text-dominant Issue. We analyze how certain pre-training strategies and model architectures exacerbate reliance on textual features during multimodal alignment. In Fig. 5(a), both PreFLMR and w/ Residual exhibit faster loss convergence compared to our model, suggesting that directly leveraging text features accelerates optimization. However, as shown in Fig. 5(b), the accelerated convergence does not translate to improved performance, with PreFLMR and w/ Residual underperforming relative to our model. The text-dominant issue is further exacerbated when using VL-ICT, introduced in ReViz, which constructs pseudo-queries from passages. Such behavior reveals the text-dominant issue, where excessive dependence on text features during alignment hinders the model’s ability to fully leverage multimodal information. Fig. 6 illustrates this effect by visualizing the alignment of multimodal query embeddings $Q$ with passage embeddings $D$. In (a), when $E_{t}$ is explicitly used during alignment $Q$ embeddings (orange) remain largely separated from the passage space, indicating poor alignment. In contrast, (b) demonstrates that our method effectively incorporates $Q$ embeddings (green) into the linguistic space, improving alignment. These results suggest that excessive reliance on text features inhibits the multimodal query embeddings from adapting properly to the passage space.
@@ 7 Limitations
8 Acknowledgment
This work was supported by the Agency For Defense Development by the Korean Government(UI247035TF)
@@ A.2 Details for Our Dataset
A.3 Architectural Differences: MIRe vs. PreFLMR
Our architecture differs from PreFLMR in several important aspects. In PreFLMR, visual tokens serve as queries, and the hidden states of the text encoder act as keys and values within a cross-attention mechanism. Attention scores $\mathcal{A}\in\mathbb{R}^{H\times Q\times K}$ are computed using these components, and value vectors $\mathcal{V}$ are derived from the truncated text encoder hidden states, which may cause information loss. The resulting outputs have dimensions $[H,Q,D]$ and are reshaped to $[Q,H\cdot D]$ before undergoing further reduction by the ColBERT head.
In contrast, MIRe uses textual embeddings $E_{t}$ from the ColBERT head as queries, with visual tokens split into multiple heads for keys and values. Here, $Q$ corresponds to the number of textual tokens and $K$ to the number of visual tokens, reversing the modality roles compared to PreFLMR. Value vectors are obtained directly from the visual embeddings without truncation. Output aggregation is performed by mean-pooling across the sequence dimension, producing representations of shape $[H,D]$, which are passed through a linear layer without further dimensionality reduction.